8. What is the pH of a solution of 0.05M NaCN.
Answer #1

NaCN is salt of weak acid HCN and strong base NaOH

for such salts

pH = 1/2 [pKw + pKa + log C]

pKa of HCN = 9.2

pH = 1/2 [ 14 + 9.2 + log 0.05]

pH = 1/2 [21.90]

pH = 10.95